# Generative AI Beyond LLMs: System Implications of Multi-Modal Generation

### v1

@@ VI-B Trends in Temporal Attention
Figure 12: Cache hit ratios for L1 and L2 cache during Attention. Temporal Attention has 10x lower L1 cache hit rate for gemm and softmax kernels as compared to Spatial Attention.
Figure 13: Benchmark illustrating how Temporal Attention FLOPs scale exponentially with number of frames as opposed to Spatial Attention, which scales linearly.
Given that Temporal Attention is a system performance bottleneck, we create a benchmark based on [40] to project future compute requirements. Figure 13 sweeps the number of frames (x-axis) and measures the corresponding FLOP count (y-axis), which is calculated by the two main matmul operations in Attention for simplicity. While FLOP count of Spatial Attention scales linearly with increasing number of frames, Temporal Attention FLOP count scales exponentially, due to frame count being the effective sequence length. For small frame counts, Temporal Attention is less computationally expensive than Spatial Attention. However, as the number of frames increases, Temporal Attention becomes the dominating system bottleneck. Note that increasing image resolution prolongs the cross-over point, as Spatial Attention computation is highly impacted by image resolution.
@@ VII Related Work
Recent work has focused on characterizing the system characteristics of LLMs, and in particular, the transformer block. [1] analyzes the latency and model FLOPS utilization (MFU) to understand how to do efficient transformer inference. [41] analyzes the heterogeneity of GEMM computations in BERT. Given that Attention is often seen as a system bottleneck, much effort has been dedicated to reducing the amount of time spent on Attention. While many works introduce approximate Attention algorithms [42, 43, 44, 45], Flash Attention [8] proposes a tiling technique to reduce HBM accesses and thus improve the Attention bottleneck. Flash Attention V2 [46] and Flash Decoding [47] further introduce more parallelism and inference optimizations, respectively. However, while previous work performs analysis on Attention speedups and end-to-end speedups, we bring this into context of an operator breakdown to determine how system performance bottlenecks shift. Additionally, we broaden evaluations to the emerging class of TTI/TTV workloads.
Another class of work has focused on optimizing TTI/TTV models from an algorithm perspective. [48] provides an overview of recent progress in diffusion model research, and focuses on evaluating the computational efficiency of these models. Ongoing innovations in TTI/TTV model architectures introduce techniques such as mixture-of-experts to multi-modal generation [49]. [50] proposes a Multi-Architecture Multi-Expert diffusion model in order to better tailor different steps of the diffusion process to their functionality. Others propose using techniques such as Retrieval Augment Generation (RAG) to supplement model architectures [30].
@@ VIII Conclusion
In this work, we present a detailed system characterization of an emerging class of multi-modal workloads. We find that Diffusion-based TTI models exhibit unique system bottlenecks such as Convolution after Flash Attention is applied. We additionally find that unlike LLMs, sequence length is highly variable over the course of inference for Diffusion models, complicating the need to design systems for optimal sequence lengths rather than the largest sequence length possible. Finally, we investigate TTV models and find that temporal attention will likely become an increasing bottleneck as we mature in TTV generation. Through our in-depth system performance analysis, we take a critical first step towards designing efficient and deployable systems for emerging TTI/TTV workloads.
IX Acknowledgements
This research is not possible without the following colleagues at Meta. We would like to thank Uriel Singer, Adam Polyak, Yipin Zhou for their help in understanding the model requirements for text-to-image and large language models and the code base, Driss Guessous for his help on understanding PyTorch’s Scaled Dot Product Attention implementation, and Henry Estela for his support with the training infrastructure.

### v2

@@ VI-B Trends in Temporal Attention
Figure 13: Performance counters Misses Per Thousand Instructions (MPKI) for L1 and L2 Cache, Average Memory Access Time (AMAT), and Total Memory BW across Temporal and Spatial Attention kernels. Temporal Attention has 1.2-1.96x higher L2 MPKI than Spatial Attention for elementwise and gemm kernels.
Figure 14: Benchmark illustrating how Temporal Attention FLOPs scale quadratically with number of frames as opposed to Spatial Attention, which scales linearly.
Given that Temporal Attention is a system performance bottleneck, we create a benchmark based on [40] to project future compute requirements. Figure 14 sweeps the number of frames (x-axis) and measures the corresponding FLOP count (y-axis), which is calculated by the two main matmul operations in Attention for simplicity. While FLOP count of Spatial Attention scales linearly with increasing number of frames, Temporal Attention FLOP count scales in polynomial time with equation $y=AX^{2}$, due to frame count being the effective sequence length. For small frame counts, Temporal Attention is less computationally expensive than Spatial Attention. However, as the number of frames increases, Temporal Attention becomes the dominating system bottleneck. Note that increasing image resolution prolongs the cross-over point, as Spatial Attention computation is highly impacted by image resolution.
@@ VII Related Work
Recent work has focused on characterizing the system characteristics of LLMs, and in particular, the transformer block. [1] analyzes the latency and model FLOPS utilization (MFU) to understand how to do efficient transformer inference. [41] analyzes the heterogeneity of GEMM computations in BERT. Flash Attention [8] proposes a tiling technique to reduce HBM accesses and thus improve the Attention bottleneck. Flash Attention V2 [42] and Flash Decoding [43] futher introduce more parallelism and inference optimizations, respectively.
Another class of work has focused on optimizing TTI/TTV models from an algorithm perspective. [3] provides an overview of recent progress in diffusion model research, and focuses on evaluating the computational efficiency of these models. [4] proposes a Multi-Architecture Multi-Expert diffusion model in order to better tailor different steps of the diffusion process to their functionality. Others propose using Retrieval Augment Generation (RAG) techniques to supplement the model architectures [30].
@@ VIII Conclusion
In this work, we present a detailed system characterization of an emerging class of multi-modal workloads. We find that Diffusion-based TTI models exhibit unique system bottlenecks such as Convolution after Flash Attention is applied. We additionally find that unlike LLMs, sequence length changes throughout the course of inference for Diffusion models, complicating the need to design systems for optimal sequence lengths rather than the largest sequence length possible. Finally, we investigate TTV models and find that temporal attention will likely become an increasing bottleneck as we mature in TTV generation. Through our in-depth system performance analysis, we take a critical first step towards designing efficient and deployable systems for emerging TTI/TTV workloads.
Acknowledgements
This research is not possible without the following colleagues at Meta. We would like to thank Uriel Singer, Adam Polyak, Yipin Zhou to understand the model requirements for text-to-image and large language models and the code base, Driss Guessous for PyTorch’s Scaled Dot Product Attention implementation, and Henry Estela for his support with the training infrastructure.